OpenStudy (anonymous):

what is the solution of 3/5=x/15

OpenStudy (anonymous):

x = .6 * 15 x = 9 I got .6 by converting 3/5 into decimal form

OpenStudy (anonymous):

3/5=x/15 by cross multiplying 3x15=45 x x 5=5x 45/5x= 9 x=9

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) You want to start off by seeing if there is any way to simplify your fractions. \[\frac{3}{5} = \frac{x}{15} \] Multiply the left fraction be 3 \[\frac{9}{15} = \frac{x}{15} \] Your 16s cancel out! You're left with \[9 = x \rightarrow x = 9\]